The solution of the equation 3x + 7 = –20 is __________.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The solution of the equation 3x + 7 =–20 is -9.
Consider the equation
3x + 7 =–20

Transposing 7 from left hand side to right hand side it becomes -7,
3x = - 20–7

3x = - 27
x = -27/3
x=-9

Thanks